OpenStudy (anonymous):

pete makes half as much as his sister Melony. Pete makes $9.50 per hour plus an additional $40 for side work. Melony makes $11 per hour. If both of their paychecks are equal, how many hours do they work?

OpenStudy (anonymous):

what class is this?

OpenStudy (anonymous):

its called math ,algebra ..

OpenStudy (anonymous):

ok, but it sounds like you haven't done problems with 2 equations and 2 unknown variables at a time yet. That's fine... you need expressions for how much each person makes...

OpenStudy (anonymous):

The way I read it, they both work the same number of hours. H is the number of hours Pete makes 9.50 per hour + 40 for some other work. P = 9.50(H) + 40 Melony makes 11 per hour M = 11(H) And you know that their paychecks are equal, so P = M, and we can set those expressions equal. 9.50(H) + 40 = 11(H) Then collect the H terms on the right: 40 = 1.50(H) And then solve for H: 40/1.5 = H H = 26.6667 hours, or 26 hours 40 minutes. So, Pete worked 26.667 hours and made 9.50(26.667) + 40 = $293.33 and Melony worked 26.66667 hours and made 11(26.667) = $293.33

OpenStudy (anonymous):

The problem is that this question says two contradictory things... One part says Pete only makes half as much as Melony. But the end statement says that their paychecks are equal. It is hard for me to understand how both can be true... if Pete makes half as much as Melony, then their paychecks CAN'T be equal. Weird :(